### v1

# Operator Learning with Domain Decomposition for Geometry Generalization in PDE Solving

Abstract
Neural operators have become increasingly popular in solving partial differential equations (PDEs) due to their superior capability to capture intricate mappings between function spaces over complex domains. However, the data-hungry nature of operator learning inevitably poses a bottleneck for their widespread applications. At the core of the challenge lies the absence of transferability of neural operators to new geometries. To tackle this issue, we propose operator learning with domain decomposition, a local-to-global framework to solve PDEs on arbitrary geometries. Under this framework, we devise an iterative scheme Schwarz Neural Inference (SNI). This scheme allows for partitioning of the problem domain into smaller subdomains, on which local problems can be solved with neural operators, and stitching local solutions to construct a global solution. Additionally, we provide a theoretical analysis of the convergence rate and error bound. We conduct extensive experiments on several representative PDEs with diverse boundary conditions and achieve remarkable geometry generalization compared to alternative methods. These analysis and experiments demonstrate the proposed framework’s potential in addressing challenges related to geometry generalization and data efficiency.
1 Introduction
Partial differential equation (PDE) solving is of paramount importance in comprehending natural phenomena, optimizing engineering systems, and enabling multidisciplinary applications Evans, (2022). The computational cost associated with traditional PDE solvers Liu and Quek, (2013); Lu et al., (2019) has prompted the exploration of learning-based methods as potential alternatives to overcome these limitations. Neural operators Li et al., 2020b ; Li et al., (2023, 2024); Liu et al., (2023); Hao et al., (2023), as an extension of traditional neural networks, aim to learn mappings between the functional dependencies of PDEs and their corresponding solution spaces. They offer highly accurate approximations to classical numerical PDE solvers while significantly improving computational efficiency. Despite its success, operator learning, as a data-driven approach, encounters the inherent ‘chicken-and-egg’ problem, revealing an interdependence between operator learning and the availability of data. This dilemma arises from the challenge of simultaneously addressing the inefficiency of classical solvers and acquiring an ample amount of data for neural operator training.
Existing works in alleviating the above challenge explore methods for data augmentation. For instance, Lie point symmetry data augmentation (LPSDA) Brandstetter et al., (2022) generates potentially infinitely many new solutions from an existing solution by exploiting symmetries of the differential operator. However, data augmentation techniques only alleviate the data efficiency issue by increasing the diversity and size of the dataset, with the nature of generalization in PDE solving unscrutinized. In this context, we introduce the ‘ladder of generalizability’, which outlines a hierarchical structure comprising three increasingly challenging problem types: (1) coefficient, involving variations in coefficients, source term and boundary conditions while keeping the PDE and domain geometry fixed; (2) geometry, involving changes in the domain geometry while keeping the PDE fixed; and (3) PDE, encompassing the ability to handle various types of PDEs with diverse dependencies. While existing neural operators have shown capabilities in handling diverse geometries through approaches such as geometry parametrization Li et al., (2023) or coordinate representation Hao et al., (2023), they lack the ability to generalize to entirely novel geometries that differ significantly from those present in the training data distribution. The inability to quickly adapt neural operators to unseen geometries without further generating new data hinders the applicability of neural operator learning to real-world problems in industry.
To tackle this challenge, a natural idea is to break down a domain into some basic shapes where neural operator can generalize well. Domain decomposition methods (DDMs) Toselli and Widlund, (2004); Mathew, (2008) provide the suitable tool for this purpose. Thus, we introduce a local-to-global framework to solve PDEs on arbitrary geometries by combining operator learning with DDMs. This framework consists of three parts: (1) Training data generation: creation of random basic shapes and imposition of appropriate boundary conditions on these shapes. This generated data serves as the training set for the neural operator in our framework. (2) Local operator learning: neural operator training to learn solutions on basic shapes. Data augmentation is utilized to enable the neural operator to capture the intricate details and variations within these shapes. (3) Schwarz neural inference (SNI): a three-step algorithm for inference. Firstly, the computational domain is partitioned into smaller subdomains. Then, the learned operator is applied within each subdomain to obtain the local solution. Finally, an iterative process of stitching and updating the global solution is performed using additive Schwarz methods.
By combining these three components, our local-to-global framework enables the accurate PDE solving on arbitrary geometries. The training data generation ensures the availability of diverse and representative data, the local operator learning captures local behavior effectively, and the SNI algorithm handles the partitioning and integration of solutions across the entire computational domain.
Our Contributions. We summarize our contributions below:
  • We introduce a local-to-global framework that combines operator learning with DDMs. This work represents the first attempt, to our best knowledge, in directly tackling the inherent ‘chicken-and-egg’ problem through the lens of geometry generalization for operator learning.
  • We propose a novel data generation scheme and an inference algorithm SNI.
  • We theoretically analyze the convergence and the error bound of the proposed algorithm. Through comprehensive experiments, we empirically validate the effectiveness of our framework on generalizing to new geometries.
2 Problem Formulation and Preliminaries
In this section, we provide an introduction to the problem formulation and essential background on domain decomposition methods, which will be utilized throughout the entirety of the paper.
2.1 Problem formulation
Our primary focus is on stationary problems of PDEs defined in the following form:
$\mathcal{L}(u)$ $=f\quad\quad\text{in}\ \ \Omega$ (1)
$u$ $=u_{D}\quad\ \text{on}\ \ \Gamma_{D}$
$\frac{\partial u}{\partial n}$ $=g\quad\quad\text{on}\ \ \Gamma_{N}$
where $\mathcal{L}$ is a self-adjoint and coercive elliptic partial differential operator Mathew, (2008) and $\Gamma_{D}\cup\Gamma_{N}=\partial\Omega$ denotes Dirichlet and Neumann boundary, respectively. We assume all the domains $\Omega$ are bounded orientable manifolds embedded in some ambient Euclidean space $\mathbb{R}^{n}$ Li et al., (2023). Later we will extend our method to handle time-dependent parabolic equations.
We consider situations where geometry of domain $\Omega_{\text{inf}}$ at inference time is decoupled from that of $\Omega_{\text{train}}$ in training time, i.e., $\Omega_{\text{inf}}$ does not have to fall in or resemble training geometries and can be of arbitrary shapes. For implementation we will mainly focus on $\Omega\subseteq\mathbb{R}^{2}$.
2.2 Domain decomposition methods
Domain decomposition methods (DDMs) solve Eq. (1) by decomposing domain into subdomains and iteratively solve a coupled system of equations on each subdomain. An overlapping decomposition of $\Omega$ is a collection of open subregions $\{\Omega_{k}\}_{k=1}^{K}$, $\Omega_{k}\subseteq\Omega$ for $k=1,\dots,K$ such that $\bigcup_{k=1}^{K}\Omega_{k}=\Omega$. We denote $V$ and $\{V_{k}\}_{k=1}^{K}$ to be finite element space associated with domain $\Omega$ and $\{\Omega_{k}\}_{k=1}^{K}$. We can define restriction operators $\{R_{k}:V\rightarrow V_{k}\}_{k=1}^{K}$ restricting functions on $\Omega$ to $\{\Omega_{k}\}_{k=1}^{K}$ and extension operators $\{R^{\intercal}_{k}:V_{k}\rightarrow V\}_{k=1}^{K}$ extending functions on $\{\Omega_{k}\}_{k=1}^{K}$ to $\Omega$ by zero.
In the subsequent discussion, we revisit the idea of additive Schwarz method (ASM) in DDMs for overlapping decomposition. The additive Schwarz-Richardson iteration Mathew, (2008) has the following form:
$u^{n+1}=u^{n}+\tau\sum_{k=1}^{K}\left[R_{k}^{\intercal}w_{k}^{n+1}-R_{k}^{\intercal}R_{k}u^{n}\right]$ (2)
where $0<\tau<\frac{1}{K}$ is a hyperparameter controlling the convergence rate, and $w_{k}^{n+1}$ is the solution of the following equation:
$\mathcal{L}(w_{k}^{n+1})$ $=0\quad\quad\text{in}\ \ \Omega_{k}$ (3)
$w_{k}^{n+1}$ $=u_{D}\quad\ \text{on}\ \ \partial\Omega_{k}\cap\Gamma_{D}$
$\frac{\partial w_{k}^{n+1}}{\partial n}$ $=g\quad\quad\text{on}\ \ \partial\Omega_{k}\cap\Gamma_{N}$
$w_{k}^{n+1}$ $=u^{n}\quad\ \ \text{on}\ \ \partial\Omega_{k}\cap\Omega$
We denote the local operator $\mathcal{S}_{k}:(u^{n},u_{D},g)\mapsto w_{k}^{n+1}$. Note that the first two boundary conditions in Eq. (3) is the boundary condition on the global boundary part of $\partial\Omega_{k}$ and is not updated during iteration. The last boundary condition is along the artificial boundary created by decomposition and the value is updated through iteration. Hence $\{\mathcal{S}_{k}\}_{k=1}^{K}$ can be considered as a single-input operator when the global boundary condition and decomposition are determined. This iterative process can be shown to converge for FEMs under mild assumption on properties of equation and decomposition. Please refer to Appendix A.1 for more details.
3 Operator Learning with Domain Decomposition
Figure 1: An illustration of Operator Learning with Domain Decomposition Framework. (a) During training stage, the goal is to ensure that the neural operator can effectively model the local solution operator on various building blocks of shapes. These building blocks are selected and generated based on specific criteria, allowing for a more efficient and targeted learning process. Proper boundary conditions are then imposed to generate local solutions which serve as training data for neural operator. (b) During inference, for an arbitrary given domain, an automated decomposition algorithm is employed to decompose the domain into subdomains. By leveraging the trained local operator and Schwarz Neural Inference (SNI), global solution can be obtained by stitching local solutions on subdomains.
In order to solve PDE on arbitrary geometry with neural operator, a natural idea is to decompose domain into a prescribed family of building blocks (basic shapes) since it is not feasible to explicitly consider arbitrary shapes during training process. The idea of marrying operator learning with domain decomposition is to train a neural operator to solve local problems on these shapes and stitch local solutions together to get a global solution, corresponding to the training and inference stages, respectively. An illustration of the proposed framework is presented in Figure 1. A detailed implementation will be discussed in the following subsections.
3.1 Training Data Generation
Data generation serves the purpose of operator learning, which fundamentally aims to approximate the local solution operator $\mathcal{G}:\mathcal{P}\times\mathcal{H}\to\mathcal{U}$. Here, $\mathcal{P}$ denotes the space of basic shapes, $\mathcal{H}$ represents boundary conditions and other input functions, $\mathcal{U}$ represents the solution space. Next we will delve into a comprehensive examination of how $\mathcal{P}$ and $\mathcal{H}$ are determined separately.
Choice of basic shapes. The selection of basic shapes cannot be arbitrary due to the requirement of ensuring the neural operator’s capability in solving local problems across a wide range of shapes. To address this issue, we need to specify a probability space $(\mathcal{P},\mu)$ where $\mu$ denote the probability distribution over $\mathcal{P}$. Moreover, two necessary criteria should be set forth for basic shape generation: (1) sampling feasibility: it should be tractable to sample from $\mu$ and solve boundary value problems on shapes in $\mathcal{P}$ in order to generate data for operator learning; (2) complete coverage: basic shapes in $\mathcal{P}$ should be flexible to cover any shape of domain.
For implementation, we focus on $\Omega\subseteq\mathbb{R}^{2}$. We propose to use the space $\mathcal{P}_{s}(n)$ of simple polygons with at most $n$ vertices (i.e. planar polygon without self-intersection and holes) uniformly bounded by a compact region in $\mathbb{R}^{2}$. Simple polygons are Lipschitz domains with straightforward sampling method Auery and Heldz, (2019) and flexible enough to constitute any discretized planar domain Preparata and Shamos, (2012). We note, however, that this is not the only choice of these basic shapes. We could equally use convex polygons, star-shaped polygons, etc. as long as the two aforementioned criteria are satisfied.
Imposing boundary conditions. The imposition of boundary conditions presents two complications: (1) Types of boundary conditions. Neumann boundary conditions in Eq. (1) will inevitably result in mixed boundary conditions in local subdomains. To generate solutions with mixed boundary conditions, we randomly divide the boundary of a basic shape into two connected components, representing the Dirichlet and Neumann boundaries, respectively. During inference, we have to carefully set hyperparameters for decomposition to make sure boundary of subdomains have at most two connected components for Dirichlet and Neumann boundaries. (2) Functional range of boundary conditions. In general, the inference process for subdomains will encounter arbitrary ranges in boundary conditions. However, it is practically infeasible to train the neural operator to handle unbounded boundary values. Instead, we generate random functions with values normalized within a bounded range for both boundary conditions and other input functions such as coefficient fields and source terms. We will handle this complication with symmetries of PDEs during inference.
3.2 Local Operator Learning
We now train a neural operator $\mathcal{G}^{\dagger}$ to approximate the mapping $\mathcal{G}$. Our focus is not on design of neural operators, but on ensuring that the neural operator can solve local problems accurately.
Choice of neural operator architecture. Our framework is orthogonal to the choice of neural operator architecture as long as the architecture can accommodate flexible input/output formats and possesses sufficient expressive power to solve local problem with randomly varying domain and input functions. For implementation, we adopt GNOT Hao et al., (2023) which is a highly flexible transformer-based neural operator. We note that, however, training neural operator on highly varying geometries presents challenges to both design of architectures and training schemes.
Data augmentation. To enhance the generalization capabilities of the neural operator, Lie point symmetry data augmentation (LPSDA) Brandstetter et al., (2022) can be naturally applied to local solutions during training. Examples of such transformations are rotation and scaling. It is crucial to appropriately extend these transformations to boundary conditions and other input functions, taking into account the symmetries inherent in the PDEs. Please refer to Appendix A.5 for a detailed discussion.
3.3 Schwarz Neural Inference
Inspired by additive Schwarz method, we introduce a similar iterative algorithm called Schwarz Neural Inference (SNI), which is outlined in Algorithm 1. In the subsequent discussion, we will explore several important considerations.
Algorithm 1 Schwarz Neural Inference
0:  Domain $\Omega$; Global Boundary Condition $B$; Other input functions $H$; Number of Subdomains $K$; Depth of Extension $d$; Local Operator $\mathcal{G}^{\dagger}$; Step Size $\tau$; Convergence Criterion $C$;
0:  Global Solution $u$;
1:  Apply METIS and extension to get overlapping decomposition $\{\Omega_{k}\}_{k=1}^{K}$, obtain restriction operators $\{R_{k}\}_{k=1}^{K}$ and extension operators $\{R_{k}^{\intercal}\}_{k=1}^{K}$;
2:  Initialize the global solution $u^{0}$;
3:  while convergence criterion $C$ not satisfied do
4:     update local boundary condition $\{B_{k}^{n}\}_{k=1}^{K}$ by global boundary condition $B$ and last-step global solution $u^{n}$;
5:     obtain the preprocessing $\{T_{k}\}_{k=1}^{K}$ and postprocessing transformations $\{\tilde{T}_{k}\}_{k=1}^{K}$;
6:     inference on each subdomain using local operator: $\tilde{w}^{n+1}_{k}=\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}(\Omega_{k},B_{k}^{n})$;
7:     extend local solution: $w^{n+1}=\sum_{k=1}^{K}R_{k}^{\intercal}\tilde{w}^{n+1}_{k}+(I-R_{k}^{\intercal}R_{k})u^{n}$;
8:     update global solution: $u^{n+1}=(1-\tau K)u^{n}+\tau w^{n+1}$;
9:     $n=n+1$;
10:  end while
11:  return  $u^{n}$;
Decomposition into overlapping subdomains. In general, there is no natural methods to decompose an arbitrary domain into desired shapes, and here we adopt the common practice in DDM literature Mathew, (2008). We assume there exists a pre-defined triangulation $\mathcal{T}_{h}(\Omega)$ of the domain $\Omega$, and a graph can be constructed to represent the connectivity of this triangulation. A graph partition algorithm such as METIS Karypis and Kumar, (1997) is then employed to partition this graph into $K$ non-overlapping connected subgraphs with index sets ~$\tilde{\mathcal{I}}_{1},\dots,\tilde{\mathcal{I}}_{K}$. To achieve an overlapping decomposition, each subgraph is then extended iteratively by including neighboring vertices for $d$ iterations. This process generates index sets $\mathcal{I}_{1},\dots,\mathcal{I}_{K}$ that, together with the original mesh, form an overlapping decomposition denoted as $\{\Omega_{k}\}_{k=1}^{K}$. An intuitive illustration of this process is depicted in Figure 1.
For implementation, partition number $K$ and extension depth $d$ are hyperparameters that should be carefully set to ensure that the resulting subdomains resemble shapes in $\mathcal{P}$.
Normalization. During inference on an arbitrary decomposed subdomain, the range of geometry and boundary conditions may differ from that of the generated training data. We thus leverage the symmetry properties of PDEs to handle this mismatch. More specifically, we can directly apply transformations $T:\mathcal{P}\times\mathcal{H}\to\mathcal{P}\times\mathcal{H}$ such as spatial translation and scaling laws to transform a local problem outside our training range - geometry or function values - into the training range. Note that the transformations have to be extended to any coefficient fields or source term if they are also involved in the symmetry. After neural operator inference, the resulting solution function will be transformed back by a proper inverse transformation $\tilde{T}:\mathcal{U}\to\mathcal{U}$. We implement these transformations as preprocessing and postprocessing steps in the inference pipeline.
Time Complexity. Suppose the single inference time of local operator and the number of iterations are denoted as $b$ and $N$. Let $v,e,K$ denote the number of vertices, edges and subdomains respectively. Our Algorithm 1 consists of two main parts: mesh partition using the METIS algorithm, the time complexity of which is approximately $O(v+e+K\log K)$ Karypis and Kumar, (1997); iterative scheme using the additive Schwarz method with a time complexity roughly $O(bKN)$. Therefore, the overall time complexity of our algorithm can be approximated as $O(v+e+K\log K+bKN)$ [^1] With neural operators implementing linear transformers, e.g., GNOT applied in this work, $b=O(\frac{v}{K}).$. While providing an exact time complexity analysis for FEM can be challenging due to the complexity and variability of different problem setups, it is worth noting that FEM is generally considered to be computationally demanding.
3.4 Theoretical Results
Here we provide a theoretical analysis of our proposed algorithm by stating the following result:
Theorem 1.
Let $u$ and $\tilde{u}$ denote the solution obtained by classical additive Schwarz method given Eq. (2) and SNI in Algorithm 1, respectively, with the same initial condition $u^{0}=\tilde{u}^{0}$. Assume $\parallel\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}(\cdot)-\mathcal{S}_{k}(\cdot)\parallel<c$ for all $k$, and if the classical algorithm converges, then we have:
  • Convergence: SNI converges to a fixed point;
  • Error bound: there exists a constant $c^{\prime}$ (depending on $c$) such that $\parallel\tilde{u}^{n}-u^{n}\parallel<c^{\prime}$.
The theorem suggests that if our learned local operator maintains a uniform error bound, the algorithm converges and exhibits a minimal approximation error. See Appendix A.3 for a proof.
4 Experiments
In this section, we perform comprehensive experiments to showcase the effectiveness of our method on various challenging datasets.
Figure 2: Illustration of experiment domain A, B, C from left to right respectively.
4.1 Experimental Setup
Datasets. To demonstrate the scalability and superiority of our method, we construct several datasets on multiple PDEs. We also extend our framework to a time-dependent problem, heat conduction. To aggregate training sets, we generate random simple polygons bounded by the unit square $[-0.5,0.5]^{2}\subset\mathbb{R}^{2}$. Boundary/initial conditions and coefficient functions are piecewise linear functions determined by random values within $[0,1]$. For each of the following problems, we test on datasets based on three different domains A, B and C shown in Figure 2. Details of these datasets are given in Appendix A.7.1.
  • Laplace2d-Dirichlet: Laplace equation in 2d with pure Dirichlet boundary condition on various shapes.
  • Laplace2d-Mixed: Laplace equation in 2d with mixed Dirichlet and Neumann boundary condition on various shapes.
  • Darcy2d: Darcy flow in 2d with coefficient field $a(x)$, source term $f(x)$ and pure Dirichlet boundary condition on various shapes.
  • Heat2d: Time-dependent heat equation in 2d with a coefficient $\alpha$ for thermal diffusivity, initial condition and time-varying pure Dirichlet boundary condition on various shapes.
Baseline. We compare our method with a direct inference of the trained neural operator on domains shifted and scaled to $[-0.5,0.5]^{2}$.
Evaluation Protocol. The evaluation metric we utilize is the mean $l_{2}$ relative error. See Appendix A.7.2 for details.
4.2 Main Results and Analysis
Equation Domain GNOT SNI
Laplace2d-Dirichlet A (21.53$\pm$2.46)e-2 (2.15$\pm$0.58)e-2
B (22.39$\pm$1.89)e-2 (2.10$\pm$0.40)e-2
C (28.45$\pm$2.82)e-2 (2.10$\pm$0.86)e-2
Laplace2d-Mixed A (10.74$\pm$0.79)e-2 (6.36$\pm$3.70)e-2
B (10.74$\pm$0.79)e-2 (7.00$\pm$1.32)e-2
C (37.53$\pm$5.71)e-2 (6.22$\pm$1.28)e-2
Darcy2d A (15.65$\pm$1.25)e-2 (8.60$\pm$1.61)e-2
B (63.42$\pm$3.01)e-2 (7.94$\pm$1.65)e-2
C (166.88$\pm$7.74)e-2 (5.35$\pm$0.64)e-2
Heat2d A (11.5$\pm$0.60)e-2 (5.28$\pm$0.16)e-2
B (31.1$\pm$14.8)e-2 (11.5$\pm$1.70)e-2
C (22.3$\pm$14.6)e-2 (5.76$\pm$0.27)e-2
Table 1: Main results. The $l_{2}$ relative errors along with standard deviation over different random boundary/initial conditions on three domains are reported.
The main results for all datasets are shown in Table 1. More details and hyperparameters are summarized in Appendix A.7.2 due to limited space. Based on these results, we have the following observations.
(a) Lap2d-D domain A
(b) Lap2d-D domain B
(c) Lap2d-D domain C
(d) Darcy2d domain A
(e) Darcy2d domain B
(f) Darcy2d domain C
Figure 3: Comparison between the $l_{2}$ relative errors from SNI (blue), GNOT direct inference (orange) and validation (red) on two different PDEs (Laplace2d-Dirichlet and Darcy2d) upon three domains (A, B and C) with different numbers of training samples. The results of SNI and GNOT direct inference are presented based on 100 inferences with different boundary conditions. The best validation errors during training are also provided as a reference.
Stationary Problems. First, we find that our method performs significantly better on all stationary problems compared with baseline. On all domains, we reduce prediction error by 34.8%-96.8%. The excellent performance shows the effectiveness of our framework in dealing with arbitrary geometries unseen during training. In particular, our framework usually leads by a larger margin on more complicated domain, due to the fact that simple polygons used in the training data fail to adequately resemble the complex testing domains. Solutions on multiply connected domains usually exhibit characteristics that are not present on simple domains.
Second, we find that the the performance of our method is consistent across various geometries during inference. On all types of PDEs in our datasets, the difference in prediction error over various geometries is within 3.25%, showing the ability to solve PDE with consistent accuracy on various geometries with a single trained neural operator. This also provides evidence for our theoretical result in Theorem 1 where we show that the SNI ensure the convergence to an approximation of the ground-truth solution with error bound determined by the generalization error of the neural operator.
Third, we find that complexity of the PDE together with types of boundary condition affect the generalizability of the neural operator in solving local problems and thus also the accuracy of our method. For simple problem such as Laplace2d-Dirichlet, our method achieve a 59.8% lower error compared to other problems. For Laplace2d-Mixed, neural operator struggles to capture subtlety in presence of both Dirichlet and Neumann boundaries. The complexity of Darcy2d lies in the need to capture changes in coefficient and source term in addition to geometry and boundary condition. We argue that having a strong neural operator that can generalize well on all basic shapes and boundary conditions is necessary for our framework to work with reasonable accuracy.
Time-dependent Problems. There is a natural way to extend our framework to time-dependent problems Li and Cai, (2015) where a space-time decomposition is constructed by taking the product of a spatial decomposition and a temporal decomposition. We train a neural operator that can predict heat conduction on multiple time steps and the same SNI is applied during inference on this 3d problem. Our framework works well on this problem and reduce prediction error by 54.1%-74.2%. This demonstrates the potential of our framework to handle time-dependent problems. We refer to Appendix A.4 for detailed implementation.
Data Efficiency. The exploration results on data efficiency of SNI are shown in Figure 3, implying the following observations: (1) At all abundances of data, the $l_{2}$ relative errors of SNI are significantly lower than those of GNOT direct inferences; (2) Errors of SNI are comparable to or even lower than validation errors at large data volumes. (3) SNI requires much smaller datasets to achieve comparable results to GNOT direct inference. Overall, these results demonstrate that SNI has substantial advantages in terms of data efficiency. Our proposed framework possesses remarkable ability to extract more insights from limited data and scale more effectively as data volumes increase. More supplementary results are provided in Appendix A.7.3.
4.3 Ablation Experiments
(a) $d=2$, $\tau=0.04$, varying $K$
(b) $K=20$, $\tau=0.04$, varying $d$
(c) $K=20$, $d=2$, varying $\tau$
Figure 4: Comparison between convergence rate of SNI on Laplace2d-Dirichlet domain A.
Validation Domain A Domain B Domain C
No Data Aug 3.79e-2 (4.38$\pm$1.58)e-2 (2.95$\pm$0.58)e-2 (2.75$\pm$0.95)e-2
Rotation Only 2.50e-2 (2.15$\pm$0.58)e-2 (2.10$\pm$0.40)e-2 (2.10$\pm$0.86)e-2
Rotation + Scale [0.2, 1] 5.31e-2 (3.74$\pm$1.20)e-2 (3.37$\pm$0.50)e-2 (3.41$\pm$0.43)e-2
[0.5, 1] 3.62e-2 (2.71$\pm$0.48)e-2 (3.73$\pm$0.61)e-2 (3.21$\pm$0.57)e-2
[0.8, 1] 2.86e-2 (1.87$\pm$0.43)e-2 (3.30$\pm$0.71)e-2 (2.78$\pm$0.82)e-2
Table 2: Comparison between models trained with different data augmentations for Laplace2d-Dirichlet.
Hyperparameter Exploration. The number of partitions ($K$), the depth of extension ($d$) and step size ($\tau$) are the key main hyperparameters that can affect the performance of SNI. Based on the results presented in Figure 4, the factors analyzed have no significant impact on the accuracy of our algorithm, but they do influence the convergence rate. Specifically, increasing the number of partitions leads to a smaller $l_{2}$ relative error but slower convergence. Once the partition number surpasses 20, the algorithm’s final performances become comparable. However, when using 50 partitions, the inference process does not converge in 5000 iterations. Regarding the depth of extension, it does not affect the performance on the tested domains. The convergence curves for depth of extension 1, 2, 4, and 8 are nearly identical. When it comes to $\tau$, a larger value results in faster convergence. However, it is important to note that there exists a maximum limit $1/K$ beyond which $\tau$ cannot be set.
Data Augmentation Exploration. To explore the effects of data augmentation, we compare the performances of models trained with different degrees of data augmentation for Laplace2d-Dirichlet demonstrated in Table 2. For models trained without data augmentation, the variance of performances on domains is large, ranging from $2.75\%$ to $4.38\%$. Specifically, it reports a $4.38\pm 1.58\%$ $l_{2}$ relative error on domain A, while this error can be reduced to $1.87\pm 0.43\%$ with a rotation+[0.8, 1] scaling augmentation. While rotation can generally be beneficial, the effectiveness of scaling can sometimes be limited or even detrimental. Hence, it is important to apply data augmentation with caution and consider its suitability for different types of PDEs.
5 Related Work
Operator Learning. The idea of operator learning is first introduced in Lu et al., (2019). This work proposes a notable architecture called DeepONet, which employs a branch network for processing input functions and a trunk network for handling query points. Adopting the trunk-branch architecture and utilizing the attention mechanism,  Hao et al., (2023) develops GNOT to handle irregular mesh, multiple input functions, and different input data types. The high accuracy and versatility makes GNOT the benchmark in our work. In the other direction, Fourier neural operator (FNO) Li et al., 2020b leverages the Fast Fourier Transform (FFT) to learn operators in the spectral domain, and achieves a favorable trade-off between cost and accuracy. Variants of FNO are proposed to reduce computational cost (FFNO in Tran et al., (2021)), handle irregular mesh (GeoFNO in Li et al., (2023)), and improve expressivity (UFNO in Wen et al., (2022)).
Domain Decomposition Methods Applied in Deep Learning. In general, the integration of deep learning and DDMs can be categorized into two groups Heinlein et al., (2021). The first category involves using deep learning techniques to improve the convergence properties or computational efficiency of DDMs. For instance, several methods Heinlein et al., (2020, 2019) have been proposed to reduce the computational cost in adaptive FETI-DP solvers by incorporating deep neural networks (DNNs) while ensuring the robustness and convergence behavior. The second category is centered around the substitution of subdomain solvers in DDMs with neural networks. There have been multiple endeavors to employ PINNs or Deep Ritz methods as alternatives to subdomain solvers or discretization techniques in traditional DDMs Li et al., 2020a ; Li et al., (2019). These approaches leverage the universal approximation capabilities of neural networks to represent solutions of PDEs, subject to specific assumptions regarding the activation function and other factors.
Data Augmentation Techniques in Operator Learning. Different types of data augmentations are proposed to improve the generalization capabilities in operator learning. A Lie point symmetry framework is introduced in Brandstetter et al., (2022), which quantitatively derives a comprehensive set of data transformations, to reduce the sample complexity. Motivated by this approach,  Mialon et al., (2023) learn general-purpose representations of PDEs from heterogeneous data by implementing joint embedding methods for self-supervised learning. An alternative research approach Fanaskov et al., (2023) introduces a computationally efficient augmentation strategy that relies on general covariance and straightforward random coordinate transformations. In general, applying data augmentation techniques for PDE operator learning can be challenging due to the unique nature of PDE theory.
6 Conclusion and Future Works
We presented a local-to-global framework based on DDMs to address the geometry generalization and data efficiency issue in operator learning. Our framework includes a novel data generation scheme and an iterative inference algorithm SNI. Additionally, we provided a theoretical analysis of the convergence and error bound of the algorithm. We conducted extensive experiments to demonstrate the effectiveness of our framework and validate our theoretical result. For future works, the rich literature of DDMs when combined with operator learning provides many potential directions to handle higher-dimensional problems, non-overlapping decomposition and more challenging types of equations.
Appendix A Appendix / supplemental material
A.1 Background on Domain Decomposition
Domain decomposition is a widely used technique in computational science and engineering that enables the efficient solution of large-scale problems by dividing the computational domain into smaller subdomains. This approach is particularly beneficial when dealing with complex problems that cannot be solved using a single computational resource. The main idea behind domain decomposition is to break down a large computational domain into smaller, more manageable subdomains. These subdomains can be arranged in a variety of ways, such as overlapping or non-overlapping, depending on the specific problem and the desired computational approach.
In this work, we decompose our domain into subdomains and adopt the hybrid formulation of Eq. (1) following (Mathew,, 2008, Section 1.1). A decomposition of $\Omega$ is a collection of open subregions $\{\Omega_{k}\}_{k=1}^{K}$, $\Omega_{k}\subseteq\Omega$ for $k=1,\dots,K$ such that $\bigcup_{k=1}^{K}\overline{\Omega_{k}}=\overline{\Omega}$. This decomposition is referred to as non-overlapping if in addition, $\Omega_{i}\cap\Omega_{j}=\varnothing$ for any $i\neq j$. Alternatively, an overlapping decomposition is one satisfying $\bigcup_{k=1}^{K}\Omega_{k}=\Omega$. Typically, a non-overlapping decomposition is one where subdomains do not intersect with each other in the interior while an overlapping decomposition constructed in practice has overlapping neighboring subdomains.
Given a decomposition of $\Omega$, a hybrid formulation of Eq. (1) is a coupled system of local PDEs on subdomains $\Omega_{k}$ equivalent to Eq. (1) satisfying two requirements. First, the restriction $u_{k}(x)$ of the solution $u(x)$ of Eq. (1) to each domain $\Omega_{k}$ must solve the local PDE, thus ensures that the hybrid formulation is consistent with the original problem in Eq. (1). Second, the hybrid formulation must be well posed as a coupled system of PDEs in the sense of Evans, (2022), i.e. its solution must exist, be unique and depend continuously on given input function and boundary/initial conditions. Intuitively, a hybrid formulation consists of a local problem posed on each subdomain and matching conditions that couples the local problems.
In this work we focus on the earliest and most elementary formulation termed Schwarz hybrid formulation (Mathew,, 2008, 1.2) based on overlapping decomposition and is applicable to a wide class of self-adjoint and coercive elliptic equations. Given an overlapping decomposition, $\partial\Omega_{k}$ can be decomposed into two disjoint parts. One (possibly empty) part $\Gamma_{k}=\partial\Omega_{k}\cap\partial\Omega$ is located in the boundary of $\Omega$ and the global boundary condition should be imposed. The other part $B_{k}=\partial\Omega_{k}\cap\Omega$ is a nonempty artificial boundary from the overlapping decomposition and a Dirichlet boundary condition from the coupling of local problems is imposed.
We refer to Mathew, (2008) for a strict definition. As an illustrative example, assume we have an overlapping decomposition with $K=2$ and consider as the original problem Laplace equation with mixed Dirichlet and Neumann boundary conditions. The following coupled system of two local PDEs is a Schwarz hybrid formulation of the original problem and solving the original equation is equivalent to solving this coupled system.
$\Delta u_{1}&=0\quad\quad\quad\ \text{in}\ \ \Omega_{1}\\ u_{1}&=u_{2}|_{\partial\Omega_{1}}\quad\text{on}\ \ \partial\Omega_{1}\cap\Omega\\ u_{1}&=u_{D}\quad\quad\ \ \text{on}\ \ \partial\Omega_{1}\cap\Gamma_{D}\\ \frac{\partial u_{1}}{\partial n}&=g\quad\quad\quad\ \text{on}\ \ \partial\Omega_{1}\cap\Gamma_{N}\end{aligned}\quad\text{and}\quad\begin{aligned} \Delta u_{2}&=0\quad\quad\quad\ \text{in}\ \ \Omega_{2}\\ u_{2}&=u_{1}|_{\partial\Omega_{2}}\quad\text{on}\ \ \partial\Omega_{2}\cap\Omega\\ u_{2}&=u_{D}\quad\quad\ \ \text{on}\ \ \partial\Omega_{2}\cap\Gamma_{D}\\ \frac{\partial u_{2}}{\partial n}&=g\quad\quad\quad\ \text{on}\ \ \partial\Omega_{2}\cap\Gamma_{N}\\ \end{aligned}$
Based on the Schwarz hybrid formulation, there are various iterative schemes with different parallelism and convergence rate. In the subsequent discussion, our focus is primarily on introducing the additive Schwarz methods (ASM). The ASM is a highly parallel algorithm Mathew, (2008) in solving the coupled system from Schwarz hybrid formulation. We briefly introduce ASM with finite element methods and refer to Gander et al., (2008) and Mathew, (2008) for details.
Assume that under weak formulation of Eq. (1) and finite element space $V$, Eq. (1) has the form $Au=f$ where $A$ is the stiffness matrix. Given an overlapping decomposition $\{\Omega_{i}\}_{k=1}^{K}$ compatible with the finite element space on $\Omega$, we have $V=\sum_{k=1}^{K}V_{k}$ as sum of local finite element subspaces $V_{k}$ on $\Omega_{k}$ and we can define local stiffness matrices $A_{k}:V_{k}\to V_{k}$, restriction operators $\{R_{k}\}_{k=1}^{K}$ restricting $V$ to $V_{k}$ and extension operators $\{R_{k}^{\intercal}\}_{k=1}^{K}$ extending $V_{k}$ to $V$ by zeros extension. We then define operators $P_{k}:V\to V$ by $P_{k}=R_{k}^{\intercal}A_{k}^{-1}R_{k}A$. Additive Schwarz operator is then defined as the sum $P_{\text{ad}}=\sum_{k=1}^{K}P_{k}$. This operator can be show to be self-adjoint and coercive and we have the following equivalence.
$Au=f\iff P_{\text{ad}}u=\sum_{k=1}^{K}R_{k}^{\intercal}A_{k}^{-1}R_{k}f$ (4)
We note that the right hand side of Eq. (4) is a preconditioned version of the left hand side. The Richardson iteration for this preconditioned problem has the following form.
$u^{n+1}=u^{n}+\tau\sum_{k=1}^{K}R_{k}^{\intercal}A_{k}^{-1}R_{k}(f-Au^{n})$ (5)
In the composite operator $R_{k}^{\intercal}A_{k}^{-1}R_{k}$, the operator $R_{k}$ first restrict a function to $\Omega_{k}$, $A_{k}^{-1}$ solve the local problem and $R_{k}^{\intercal}$ extend the local solution to $\Omega$. This iterative process can be shown to converge by estimating bound on condition number of $P_{\text{ad}}$ under mild assumptions on equation and decomposition.
A.2 Revisit on Operator Learning
The goal of operator learning is to learn a mapping $\mathcal{G}:\mathcal{A}\to\mathcal{U}$ between two infinitely dimensional spaces Kovachki et al., (2023). When applied to PDEs, $\mathcal{U}$ is the solution space of a PDE and $\mathcal{A}$ is the space of functions that determine a unique solution of a PDE. Examples of $\mathcal{A}$ are coefficient functions or boundary/initial conditions that defines the PDE and parameters that determine the geometry of domain.
In our study, we decompose any domain into subdomains each of which lives in a distinguished class of basic shapes $\mathcal{P}$. We assume all shapes in $\mathcal{P}$ have Lipschitz boundary and are uniformly bounded, i.e., they are all bounded by a ball $D\subseteq\mathbb{R}^{n}$. We are interested in solving boundary value problems in Eq. (1) in any domain $\Omega\in\mathcal{P}$ with any appropriate boundary condition. We thus separate geometry and boundary conditions from other inputs and represent the input function space of the operator as $\mathcal{A}=\mathcal{P}\times H^{k}(D)\times\mathcal{H}$ where $H^{k}(D)$ is the Sobolev space $W^{k,2}(D)$. The space $\mathcal{P}\times H^{k}(D)$ represents the geometry of the domain together with boundary/initial conditions, $\mathcal{H}$ represents any other input functions such as coefficient function field or source term in the PDE. The neural operator thus approximates the following mapping. Note that in the case of time dependent problem, the space $H^{k}(D;\mathbb{R})$ represents the space of initial condition together with time varying boundary condition and the solution space $\mathcal{U}$ represents a time series up to some time span. The solution operator $\mathcal{G}$ thus has the following form.
$\mathcal{G}:\mathcal{P}\times H^{k}(D)\times\mathcal{H}\to\mathcal{U}$ (6)
For learning the operator, we assume $\mathcal{P}$, $H^{k}(D)$ and $\mathcal{H}$ are probability spaces and thus we can sample observations from $\mathcal{A}$. In practice, we randomly sample geometry from $\mathcal{P}$ and random boundary conditions are imposed, then a solution is generated from a numerical solver to get solutions. It is important to highlight that, unlike the usual setting for neural operators, there is significant variation in the shape of input domains.
A.3 Proof of Theorem 1
Theorem.
Let $u$ and $\tilde{u}$ denote the solution obtained by traditional additive Schwarz method given by Eq. (5) and SNI in Algorithm 1, respectively, with the same initial condition $u^{0}=\tilde{u}^{0}$. Assume $\parallel\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}(\cdot)-\mathcal{S}_{k}(\cdot)\parallel<c$ where $\mathcal{S}_{k}$ is the local solution operator for Eq. 3 for all $k$, and if the traditional algorithm converges, then we have:
  • Convergence: SNI converges to a fixed point;
  • Error bound: there exists a constant $c^{\prime}$ (depending on $c$) such that $\parallel\tilde{u}^{n}-u^{n}\parallel<c^{\prime}$.
Proof.
(1) Recall that the iterative rules of traditional and neural Schwarz inference are given by:
$u^{n+1}$ $=u^{n}+\tau\sum_{k=1}^{K}\left[R_{k}^{\intercal}\mathcal{S}_{k}(u^{n})-R_{k}^{\intercal}R_{k}u^{n}\right]$ (7)
$\tilde{u}^{n+1}$ $=\tilde{u}^{n}+\tau\sum_{k=1}^{K}\left[R_{k}^{\intercal}\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-R_{k}^{\intercal}R_{k}\tilde{u}^{n}\right]$
where $\tilde{\mathcal{G}}_{k}^{\dagger}$ denotes $\tilde{T}_{k}\circ\mathcal{G}^{\dagger}\circ T_{k}$. By simple calculation, we first express the operator $\mathcal{S}_{k}$ explicitly by introducing finite element space under weak formulation: $\mathcal{S}_{k}:u^{n}\mapsto A_{k}^{-1}R_{k}(f-Au^{n})+R_{k}u^{n}$. Hence the convergence of traditional Schwarz algorithm implies $\rho(I-\tau MA)<1$ where $M=\sum_{k=1}^{K}R_{k}^{\intercal}\tilde{A}_{k}^{-1}R_{k}$, and $\rho(\cdot)$ denotes the spectral radius. Hence we have
$\tilde{u}^{n+1}=(I-\tau MA)\tilde{u}^{n}+\tau Mf+\tau\sum_{k=1}^{K}R_{k}^{\intercal}(\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n}))$
Since $\tau Mf$ is a constant term and $\parallel\tilde{\mathcal{G}}_{k}^{\dagger}(\cdot)-\mathcal{S}_{k}(\cdot)\parallel<c$, by random fixed point theorem, the neural algorithm converges to a random fixed point.
(2) Subtraction of Eq. (7) gives:
$\left\|\tilde{u}^{n+1}-{u}^{n+1}\right\|$ $=\left\|(I-\tau\sum_{k=1}^{K}R_{k}^{\intercal}R^{k})(\tilde{u}^{n}-{u}^{n})-\tau\sum_{k=1}^{K}R_{k}^{\intercal}(\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(u^{n}))\right\|$
$=\left\|(I-\tau\sum_{k=1}^{K}R_{k}^{\intercal}R^{k})(\tilde{u}^{n}-{u}^{n})-\tau\sum_{k=1}^{K}R_{k}^{\intercal}(\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n})+\mathcal{S}_{k}(\tilde{u}^{n})-\mathcal{S}_{k}(u^{n}))\right\|$
$\leq(I-\tau MA)\left\|\tilde{u}^{n}-{u}^{n}\right\|+\tau\sum_{k=1}^{K}R_{k}^{\intercal}\left\|\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n})\right\|$
It is easy to see that $\tau\sum_{k=1}^{K}R_{k}^{\intercal}\left\|\tilde{\mathcal{G}}_{k}^{\dagger}(\tilde{u}^{n})-\mathcal{S}_{k}(\tilde{u}^{n})\right\|\leq\tau tc$ where $t$ denotes the maximal number of overlapping subdomains. Let $\rho=\rho(I-\tau MA)<1$, we have $\left\|\tilde{u}^{n}-{u}^{n}\right\|\leq\frac{1-\rho^{n}}{1-\rho}\tau tc$. Taking $c^{\prime}=\frac{\tau tc}{1-\rho}$ completes the proof. ∎
If we apply matrix form of neural operator, namely, the neural operator aims to approximate $\{A_{k}^{-1}\}_{k=1}^{K}$ and assume, then we can have the following result:
Corollary 1.
Consider the exact operator $A_{k}^{-1}$ and inexact neural operator $\tilde{A}_{k}^{-1}$, $k=1,\cdots,K$. Let $u^{n}$ and $\tilde{u}^{n}$ represent the solutions updated by $A_{k}^{-1}$ and $\tilde{A}_{k}^{-1}$ respectively at the $n$-th step, where the updating rule is given by Eq. (5) with $\tau=1$ and both sharing the same initialization. Suppose that $\parallel A_{k}^{-1}-\tilde{A}_{k}^{-1}\parallel<c$, for $k=1,\cdots,K$, and $\rho(I-MA)<1$, where $M=\sum_{k=1}^{K}R_{k}^{\intercal}\tilde{A}_{k}^{-1}R_{k}$, then we have:
  • Convergence: the algorithm converges to a fixed point;
  • Error bound: there exists a constant $c_{1}(c)$ such that $\parallel\tilde{u}^{n}-u^{n}\parallel<\frac{c_{1}}{\parallel I-MA\parallel}$;
  • Condition number: $\kappa(MA)\leq\min(t(K+1),1+\max_{k}\frac{H_{k}}{d})$,
where $t$, $K$, $H_{k}$ and $d$ denote the maximal number of overlapping subdomains, the number of subdomains, the diameter of $k$-th subdomain, and the number of extensions, respectively.
Note that the condition $\rho(I-MA)<1$ is generally challenging to satisfy. To address this issue, we employ the Richardson iteration trick Richardson, (1911) in order to ensure the convergence of the proposed algorithm (Algorithm 1).
A.4 Time-Dependent Problems.
We consider the time-dependent PDE with the following form:
$u_{t}-\mathcal{L}u$ $=f\quad\quad\quad\quad\text{in}\ \ \Omega\times[0,T]$ (8)
$u(x,t)$ $=u_{D}(x,t)\quad\ \text{on}\ \ \partial\Omega\times[0,T]$
$u(x,0)$ $=u_{0}(x)\quad\quad\ \text{on}\ \ \Omega\times\{0\}$
where $\mathcal{L}$ is again self-adjoint and coercive elliptic operator. The additive Schwarz method can be naturally extended to a space-time additive Schwarz method Li and Cai, (2015) by considering a decomposition of the space-time domain $\Omega\times[0,T]$ by taking the product of overlapping decomposition of $\Omega$ and $[0,T]$ respectively. The space-time domain decomposition has the form $\Omega_{i}\times[t_{j-1}-\delta_{T},t_{j}+\delta_{T}]$ where $\delta_{T}$ is the temporal depth and represent overlap in time domain. Once such a decomposition is constructed, the same additive Schwarz method can be applied to the space-time decomposition to get a global solution on the space-time domain, allowing parallel iteration in both space and time domain. Local problems for the above decomposition are again of the form in Eq. (8).
In our implementation on heat equation, we discretize the time domain with a fixed time step $t_{s}$, fix a rollout length of $k$ and train a neural operator to map initial and boundary conditions to time series for the $k$ steps at $t=0,t_{s},\cdots,(k-1)t_{s}$. More precisely, the neural operator is trained to map $u_{D}$ and $u_{0}$ to time series of the form $u(x,0),u(x,t_{s}),\cdots,u(x,(k-1)t_{s})$.
A.5 Symmetries of PDEs
The symmetry group of a general partial differential operator $\mathcal{L}$ refers to a set of transformations that map a solution to another solution, forming a mathematical group. Lie point symmetry is a subgroup of the symmetry group that has a Lie group structure and acts on functions pointwise as transformations on coordinates and function values Brandstetter et al., (2022). In this work, we will in addition be concerned with not just a single operator $\mathcal{L}$, but a family of operators depending on various coefficient fields (e.g., Darcy flow) and various boundary/initial conditions. Symmetries have to be properly extended to these input functions so that a solution with an input function is transformed to another solution with a different input function.
Leveraging these symmetries allows for the generation of an infinite number of new solutions based on a given solution. The idea of utilizing these symmetries as a data augmentation technique for operator learning was initially introduced in Brandstetter et al., (2022). However, we apply these data augmentation to solutions on basic shapes in training local operator and this usage of symmetries echos a point mentioned in (Brandstetter et al.,, 2022, Section 3.2) where the authors point out that these data augmentation can be applied on local patches of solutions instead of the solution on the entire domain.
There is another direct usage of symmetries in our framework. Instead of incorporating symmetries as a form of data augmentation in training time, one can directly apply transformations to input and output of a neural operator during inference time. We implement these transformations as preprocessing and postprocessing steps in the inference pipeline. We summarize the symmetries of each PDE applied in our implementation in Table 3. Normalizations applied as preprocessing and postprocessing for each of the equations are summarized in Table 4.
Equation Lap2d-D Lap2d-M Darcy2d Heat2d
Spatial
Shift
$(x_{1},x_{2})\rightarrow(x_{1}+t_{1},x_{2}+t_{2})$
Spatial
Rotation
$(x_{1},x_{2})\rightarrow(x_{1}\cos\theta-x_{2}\sin\theta,x_{1}\sin\theta+x_{2}\cos\theta)$
Spatial Scaling $(x_{1},x_{2})\rightarrow(sx_{1},sx_{2})$
$u\rightarrow su$
$u_{D}\rightarrow su_{D}$
$g\rightarrow g$
$u\rightarrow s^{2}u$
$u_{D}\rightarrow s^{2}u_{D}$
$a(x)\rightarrow a(x)$
$f(x)\rightarrow f(x)$
$u\rightarrow u$
$u_{0}\rightarrow u_{0}$
$u_{D}\rightarrow u_{D}$
$\alpha\rightarrow s^{2}\alpha$
Value
Shift
$u\rightarrow u+t$
$u_{D}\rightarrow u_{D}+t$
Value Scaling
$u\rightarrow su$
$u_{D}\rightarrow su_{D}$
$g\rightarrow sg$ $u_{0}\rightarrow su_{0}$
Table 3: Symmetries of various PDEs applied in our implementation.
A.6 Discussions
Message passing in DDMs. In our framework, we solve a coupled system of local problems by an iterative algorithm SNI. Through iteratively solving local problems based on boundary values from the last iteration and thus from neighboring subdomains, SNI is essentially performing message passing between subdomains. This message passing operation may be implemented in other forms, e.g., through a graph neural network.
Higher-dimensional PDEs. Our framework can be extended to higher-dimensional cases as long as basic shapes and corresponding solutions can be properly generated. For 3-d problems, one potential selection of basic shapes is the class of polytopes.
Other formulations of DDMs. Schwarz hybrid formulation discussed in this work is one of the most elementary formulation in DDMs. There are many other more advanced DDMs Mathew, (2008). Steklov-Poincaré framework is based on non-overlapping decomposition and transmission condition as coupling condition for local problems. Langrange multiplier framework leads to the well-known FETI method and is also based on non-overlapping decomposition.
Other types of PDEs. The additive Schwarz method in classical DDMs works for self-adjoint and coercive elliptic equations. Non-self-adjoint elliptic equations, parabolic equations, saddle-point problems and non-linear equations requires separate treatment. Addressing these cases presents challenges in both training the local operator and designing the iterative algorithm.
Future Works. Based on the above discussion, there are many potential directions for future works. First, it would be interesting to implement this framework using a message-passing framework instead of an iterative algorithm to accelerate the convergence. Second, extending our framework to address higher-dimensional problems is important, particularly since industrial problems often involve 3-d simulation. Third, more advanced DDMs such as Neumann-Neumann, BDDC and FETI Mathew, (2008) may also be explored. Lastly, other types of PDEs such as saddle point problems and non-linear equations such as Navier-Stokes equation is out of the scope of our current work, and present unique challenge. Tackling these challenges requires not only expertise on operator learning, but also deep understanding of PDEs themselves. We speculate that it would be fruitful to combine rich literatures of DDMs with operator learning.
A.7 Experiments
A.7.1 Datasets
Figure 5: Illustration of mixed Dirichlet and Neumann boundaries for domain A, B, C in Laplace2d-Mixed.
Here we introduce more details of our datasets in both training and testing stage. For training data, we generate random simple polygons with $3\leq n\leq 12$ vertices within $[-0.5,0.5]^{2}$ and create uniform mesh using Gmsh Geuzaine and Remacle, (2008). We prepare a separate dataset for validation during training. For testing data, we generate the three domains depicted in Figure 2 together with mesh using the Gmsh UI. We argue that the complexity of geometric domains is fundamentally determined by their underlying topological and geometrical properties. Based on this intuition, we considered three domains of increasing complexity for evaluation: (1) Domain A: This domain is simply connected, representing the simplest class of geometries; (2) Domain B: This domain has two holes and is multiply connected, indicating a higher level of complexity compared to the simply connected Domain A; (3) Domain C: This domain has one hole with corners, further increasing the geometrical complexity compared to the previous two domains. Through a systematical evaluation across this spectrum of domains, from the simple geometry to more intricate multiply connected domains with holes and corners, we believe the results provide a comprehensive understanding of our framework’s capabilities.
Once the geometries and meshes are created, we specify boundary/initial conditions for various equations and domains and generate solutions using FEniCSx Baratta et al., (2023); Scroggs et al., 2022a ; Scroggs et al., 2022b , a popular open-source platform for solving PDEs with the finite element method (FEM). We adopt Lagrange element of order $1$ (linear element) as our finite element space in generating boundary/initial conditions and solutions. Next we give details on how these boundary/initial conditions and solutions are generated for each type of PDEs. We also summarize these details in Table 5 and 6.
Laplace2d-Dirichlet. Laplace equation in 2d with pure Dirichlet boundary condition. The governing equation is
$\Delta u$ $=0\quad\quad\text{in}\ \ \Omega$ (9)
$u$ $=u_{D}\quad\ \text{on}\ \ \partial\Omega.$
For both training and testing data, we specify piecewise linear Dirichlet boundary condition with randomly generated values within $[0,1]$ on boundary nodes.
Laplace2d-Mixed. Laplace equation in 2d with mixed Dirichlet and Neumann boundary condition on $\partial\Omega=\Gamma_{D}\cup\Gamma_{N}$. The governing equation is
$\Delta u$ $=0\quad\quad\text{in}\ \ \Omega$ (10)
$u$ $=u_{D}\quad\ \text{on}\ \ \Gamma_{D}$
$\frac{\partial u}{\partial n}$ $=g\quad\quad\text{on}\ \ \Gamma_{N}.$
For training data, $20\%$ of the data have pure Dirichlet condition and the generation process is the same as in Laplace2d-Dirichlet. The rest $80\%$ of the training data have mixed boundary condition with non-empty connected Neumann boundary and Neumann boundary is randomly specified to be less than half of the entire boundary. Then a random number $r$ is sampled from $U[0.5,1]$ to specify functional range for Dirichlet and Neumann boundaries as described next. Among data with non-empty Neumann boundary, $50\%$ have $u_{D}\in[0,r]$ and $g\in[0,1]$ and the other $50\%$ have $u_{D}\in[0,1]$ and $g\in[0,r]$.
For testing data, Dirichlet and Neumann boundary is specified for each of the domain A,B and C as shown in Figure 5. Boundary conditions $u_{D}$ and $g$ are both piecewise linear with randomly generated values within $[0,1]$.
Darcy2d. Darcy flow in 2d with coefficient field $a(x)$, source term $f(x)$ and pure Dirichlet boundary condition. The governing equation is
$-\nabla(a(x)\nabla u)$ $=f\quad\quad\text{in}\ \ \Omega$ (11)
$u$ $=u_{D}\quad\ \text{on}\ \ \partial\Omega.$
For training data, Dirichlet boundary condition is specified with a random range $r\in[0.3,1]$ and boundary values are generated as $u_{D}\in[0,r]$. The coefficient function $a(x)$ and source term $f(x)$ are specified as piecewise linear functions with randomly generated values within $[0,1]$ on nodes.
For testing data, Dirichlet boundary condition, coefficient function and source term are all piecewise linear functions with randomly generated values within $[0,1]$.
Heat2d. Time-dependent equation of heat conduction in 2d with coefficient $\alpha$ denoting the thermal diffusivity, time-varying boundary condition and initial condition. The governing equation is
$\frac{\partial u}{\partial t}$ $=\alpha\Delta u\quad\quad\quad\text{in}\ \ \Omega\times[0,T]$ (12)
$u(x,t)$ $=u_{D}(x,t)\quad\ \ \text{on}\ \ \partial\Omega\times[0,T]$
$u(x,0)$ $=u_{0}(x)\quad\quad\ \ \ \text{on}\ \ \Omega\times\{0\}.$
For training data, we discretize the time domain with a fixed time step $t_{s}=0.01$, generate piecewise linear initials and time-varying boundary conditions with values randomly generated within $[0,1]$. $\alpha$ is a random number within $[0.8,1]$. We adopt the backward Euler method Langtangen and Logg, (2017) and generate a time series of $10$ time steps. During training we separate these $10$ times steps into $2$ times series of $5$ times steps and training the neural operator to predict $5$ time steps.
For testing data, we fix $\alpha=1$. Initial condition is piecewise linear with values randomly generated within $[0,1]$. Boundary condition is specified to be constant over time and varied randomly within $[0,1]$ across boundary nodes. We also adopt the backward Euler method and generate a time series of $50$ time steps.
Lap2d-D Lap2d-M Darcy2d Heat2d
Data Generation Polygon [3,12]
Training
Configuration
10$\times$2000 20$\times$2000 10$\times$4000 50$\times$1600
Validation
Configuration
10$\times$250 20$\times$200 10$\times$250 50$\times$240
Testing
Configuration
100 100 100 10
Operator Learning GNOT 1 expert and 3 layers of width 128
Optimization Adam
Learning rate cycle learning rate strategy with 0.001
Epoch 500 1000 500 200
Data Aug. Rot.
Rot.+
Sca. [0.8,1]
No
Rot.+
Sca. [0.8,1]
Time steps $5$
Inference (SNI) A Partition $K$ 20 20×$\times 16$
Depth $d$ 2
Temp. Depth $\delta_{T}$ 1
Step size $\tau$ 0.04 0.002125
Pre/Post-pro.
Spa. Shift+Scale
Val. Shift+Scale
Spa. Shift+
Scale
Spa. Shift+Scale
Val. Shift+Scale
B Partition $K$ 40 40×$\times 16$
Depth $d$ 2
Temp. Depth $\delta_{T}$ 1
Step size $\tau$ 0.024 0.0014625
Pre/Post-pro.
Spa. Shift+Scale
Val. Shift+Scale
Spa. Shift+
Scale
Spa. Shift+Scale
Val. Shift+Scale
C Partition $K$ 20 20$\times$ 16
Depth $d$ 2
Temp. Depth $\delta_{T}$ 1
Step size $\tau$ 0.04 0.002125
Pre/Post-pro.
Spa. Shift+Scale
Val. Shift+Scale
Spa. Shift+
Scale
Spa. Shift+Scale
Val. Shift+Scale
Table 4: Key hyperparameters of main experiments. Configuration under Data Generation is specified as (number of random input functions per shape) $\times$ (number of various shapes). Partition $K$ for Heat2d is specified as (number of spatial partition) $\times$ (number of temporal partition).
PDE Description
Lap2d-D range of boundary condition: $U[0,1]$
Lap2d-M
20% pure Dirichlet condition: range $U[0,1]$
40% mixed boundary condition with $\Gamma_{D}/\partial\Omega\sim U[0.5,1]$
    range of Dirichlet: $U[0,r]$ where $r\sim U[0.5,1]$, range of Neumann:$U[0,1]$
40% mixed boundary condition with $\Gamma_{D}/\partial\Omega\sim U[0.5,1]$
    range of Dirichlet: $U[0,1]$, range of Neumann: $U[0,r]$ where $r\sim U[0.5,1]$
Darcy2d
range of boundary conditions: $U[0,r]$ where $r\sim U[0.3,1]$,
range of $a(x)$ and $f$: $U[0,1]$
Heat2d range of initial/boundary condition: $U[0,1]$ , $\alpha\sim U[0.8,1]$
Table 5: Details of boundary/initial condition and input function generation in training data.
PDE Description
Lap2d-D range of boundary condition: $U[0,1]$
Lap2d-M
$\Gamma_{D}$ and $\Gamma_{N}$ as in Figure 5
range of Dirichlet: $U[0,1]$, range of Neumann: $U[0,1]$
Darcy2d
range of boundary conditions: $U[0,1]$
range of $a(x)$ and $f(x)$: $U[0,1]$
Heat2d
$\alpha=1$
range of boundary/initial condition: $U[0,1]$
boundary condition do not vary with time
Table 6: Details of boundary/initial condition and input function generation in testing data.
A.7.2 Evaluation Protocol and Hyperparameters.
Evaluation Protocol. The evaluation metric we utilize is the mean $l_{2}$ relative error. Let $u_{i},u^{\prime}_{i}\in\mathbb{R}^{n}$ represent the ground truth solution and predicted solution for the $i$-th sample, respectively. Considering a dataset of size $D$, the mean $l_{2}$ relative error is computed as follows:
$\varepsilon=\frac{1}{D}\sum_{i=1}^{D}\frac{\parallel u^{\prime}_{i}-u_{i}\parallel_{2}}{\parallel u_{i}\parallel_{2}}$ (13)
Hyperparameters. All experimental hyperparameters used in the paper are listed in Table 4. For data generation, the number of vertices of simple polygons are uniformly chosen between 3 to 12. And $a\times b$ in configurations denotes the generation of $b$ shapes, each having $a$ distinct boundary/initial conditions. For investigating data efficiency issue, we only vary the number of various shapes $b$ while keeping the number of random input functions per shape $a$ constant. For boundary condition imposition, we summarize the details in Table 5 and 6.
Computing Resource. We run our experiments on 1 Tesla V100 GPU.
A.7.3 Other Supplementary Results.
Data Efficiency. The results for data efficiency on Laplace2d-Mixed problem are shown in Figure 6. The average performance of SNI is better than GNOT on all of three domains, while the margins between the two methods on domains A and B are not statistically significant due to the high variance in the $l_{2}$ relative errors of SNI on these two domains. GNOT struggles in the generalization to domain C, while SNI can still handle it with a good performance.
(a) Lap2d-M domain A
(b) Lap2d-M domain B
(c) Lap2d-M domain C
Figure 6: Comparison between the $l_{2}$ relative errors from SNI (blue), GNOT direct inference (orange) and validation (red) on Laplace2d-Mixed upon three domains (A, B and C) with different numbers of training samples.
Solution and Error Visualization. We provide visualization for stationary problems in Section 4. We visualize ground-truth solution from testing data, absolute error from SNI and GNOT direct inference in Figure 7, 8 and 9.
Figure 7: Visualization of test dataset of Laplace2d-Dirichlet on domain A, B and C. The three columns from left to right display the ground-truth solution, absolute error from SNI and GNOT direct inference.
Figure 8: Visualization of test dataset of Laplace2d-Mixed on domain A, B and C. The three columns from left to right display the ground-truth solution, absolute error from SNI and GNOT direct inference.
Figure 9: Visualization of test dataset of Darcy2d on domain A, B and C. The three columns from left to right display the ground-truth solution, absolute error from SNI and GNOT direct inference.
A.8 Broader Impacts
First, the proposed framework holds the potential to serve as an alternative to conventional PDE solving tools. Through its ability to address challenges related to geometry-generalization and data efficiency, the framework offers advantages that can significantly improve the efficiency of PDE solving. This improvement can have a positive impact on various industries, including engineering, physics, and finance, where PDEs are extensively employed for modeling and simulation purposes.
Second, the proposed three-level hierarchy for PDE generalization provides researchers with valuable directions for future exploration in neural operator research. This hierarchical structure offers a framework to systematically address the challenges associated with generalizing neural operators to new geometries and PDEs. By considering these three levels, researchers can focus on developing techniques and methodologies that improve the adaptability, flexibility, and scalability of neural operators. Furthermore, current operator learning methods in the neural operator field are predominantly driven by data and do not adequately consider the underlying PDE information. In our research, we introduce domain decomposition into the neural operator domain to tackle the issue of geometric generalization, incorporating traditional PDE approaches. This research direction presents significant potential for further investigation.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The paper’s contribution and scope are clearly stated in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitation of the paper, including future work, is presented in the A.6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The full set of assumptions are covered in the theorem, and the proof is given in A.3.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All the experimental details including hyperparameters are provided in A.7.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code will be released after the review, and we have attached our code in the supplementary files.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All the experimental details including hyperparameters are provided in A.7.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All the experimental results (Table 1, 2, Figure 3, 6) are reported with statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Computing resources for each experiment are provided in A.7.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: NeurIPS Code of Ethics is strictly followed in our research work.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Broader impacts are discussed in A.8.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper deals with simulation results of PDEs, so it poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All the creators of assets (GNOT, METIS, Schwarz, LPSDA, FEniCSx, etc.) are properly credited in our paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper deals with simulation results of PDEs, so no crowdsourcing experiments and research with human subjects is involved.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper deals with simulation results of PDEs, so no crowdsourcing experiments and research with human subjects is involved.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.